OpenStudy (anonymous):

Help!! Will give medal and fan! Which exponential function goes through the points (1, 8) and (4, 64)? f(x) = 4(2)x f(x) = 2(4)x f(x) = 4(2)-x f(x) = 2(4)-x

OpenStudy (acxbox22):

what do you think?

OpenStudy (anonymous):

4(2)^x

OpenStudy (acxbox22):

you are correct because if you plug in the x values you get the y for both points! good job

OpenStudy (anonymous):

Thanks!!!

OpenStudy (anonymous):

can I ask one more question?

OpenStudy (anonymous):

what is the product of\[4\sqrt{6}\] and\[6\sqrt{2}\]

OpenStudy (acxbox22):

multiply the outsides of the radical and then multiply the insides of the radical

OpenStudy (acxbox22):

so \[(4\times6)\sqrt{(6\times2)}\]

OpenStudy (anonymous):

48sqrt3

OpenStudy (acxbox22):

sorry, nvr mind you got it right

OpenStudy (acxbox22):

;)